7. The Petersons family has the following consumption function: C 50+2VY a. Calculate the marginal propensity to consume when γ = 36 and give an economic interpretation of your answer b. From your answer above, what is Petersons familys marginal propensity to save answer means in W
Answer #1

a. MPC+MPS =1,

We need to calculate MPC = slope = change in consumption at Y = 36,

Change in consumption = (C2-C1)/(Y2-Y1) = 0.17

Y C MPC MPS
35.00 61.83
36.00 62.00 0.17 0.83

MPC = 0.17 means, Peterson's family has 17% propensity to consume for evey 1$ income increase

b. MPS = 1-MPC = 0.83, means Peterson's family has 83 % propensity to save for evey 1$ income increase